### v2

# Conservation Laws Reveal the Quantumness of Gravity

Abstract
Adopting a general framework for quantum-classical dynamics, we analyze the interaction between quantum matter and a classical gravitational field. We point out that, assuming conservation of momentum or energy, and assuming that the dynamics obeys a particular decomposition property set out in the paper, the classical gravitational field cannot change the momentum or energy of the quantum system, whereas the quantum gravitational field can do so. Drawing upon the fundamental relationship between conservation laws and the quantum properties of objects, our analysis offers new perspectives for the study of quantum gravity and provides a novel interpretation of existing experimental observations, such as free fall.
The quantization of gravity, a fundamental issue of modern theoretical physics, has captivated the imaginations of physicists for over a century. This profound quest seeks to reconcile two pillars of contemporary physics: general relativity, which describes gravity as the curvature of spacetime, and quantum theory, which successfully accounts for the behavior of particles at the smallest scales of nature. Yet, it is hard to unify them into a unique corpus of laws. Several approaches, such as field theory, “geometrization” of quantum mechanics and string theory, have emerged in the pursuit of reconciling quantum theory with general relativity [1, 2, 3]. However, all these approaches face significant theoretical and conceptual challenges [4, 5], which have prompted some to propose the fundamental classical nature of gravity [6]. Further, some have even argued that quantizing gravity is not needed after all [7] or that gravity may not even be a fundamental force [8, 9]. Whether or not the quantisation of gravity is necessary has become highly contentious [10, 11, 12, 13, 14], primarily due to the enduring difficulty in acquiring direct supporting evidence [15, 16].
As observed by Feynman during the 1950 Chapel Hill conference, it was pointed out that the physical state of the composite system exhibits distinct properties depending on whether the mass is entangled with gravity or merely classically correlated with it [20]. This highlights the importance of the superposition of large masses in the study of quantum gravity. Recently, there has been a revival of Feynman’s idea , which highlights gravitationally induced entanglement between two massive particles provide a strong evidence on whether the gravitational field is mediated by the transfer of quantum information [18, 19]. This is clearly sufficient evidence for quantum effects in gravity. Unfortunately, gravitationally induced entanglement is extremely weak, and even though there are some advanced theoretical analyses [21, 22, 23, 24] and experimental protocols [25, 26, 27, 28, 29, 30] available, with the current experimental techniques and advances, it is not yet time to observe this effect.
The progress outlined above based on quantum information perspective is truly enlightening, even though the pursuit of quantum entanglement induced by gravity remains a formidable challenge. A natural question arises: Are the existing experimental observations of gravity sufficient to analyze the quantum nature of gravity? In this paper, we unequivocally affirm this possibility.
In historical context, Eppley and Hannah endeavored to construct a logical argument challenging the classical model of gravity’s compatibility with the conservation of momentum [10]. It is essential to acknowledge that their work predominantly consisted of a thought experiment and lacked empirical observations. Subsequently, their assertions have faced widespread skepticism and have been refuted by subsequent research [11, 13]. Moreover, other gedankenexperiments have been conducted, highlighting the perplexing paradoxes of causality and complementarity arising from classical gravity [31, 32]. Indeed, the introduction of some basic principles and physical laws helps us to analyse the non-classical nature of matter, even in the absence of entanglement witnesses [33, 34, 35, 36].
Inspired by [34, 35], our work aims to delve into the intrinsic relationship between conservation laws and the underlying nature of gravity. We shall aim at relaxing some of the assumptions in [34], e.g. linearity of the overall dynamics, by adding another constraint on the dynamics, outlined in section II. To accomplish this, we harness information-theoretic methodologies to construct a general framework of hybird systems that amalgamates quantum and classical dynamics into a unified whole [37]. Within this framework, we present a no-go theorem, establishing a fundamental limitation. It stipulates that any quantum-classical dynamics cannot induce alterations in the local observables of either classical or quantum systems, all the while maintaining the conservation of global observables. Our theorem brings to light a profound revelation: when considering the conservation of either momentum or energy, the classical gravitational field is unable to effect changes in the momentum or energy of a quantum system. This proposition stands in contrast to established experimental observations, such as the free fall experiment. Conversely, our argument compellingly demonstrates that the quantum gravitational field possesses the capability to enact such changes, offering compelling evidence for the quantum nature of gravity.
I Definition of a Classical system
A fundamental distinction between classical and quantum systems lies in their contextual nature. Classical systems operate within a single context, allowing for precise and simultaneous measurements of their dynamic variables using a single device, which can be perfectly replicated. The definition of a classical system can be established by considering some specific properties of the corresponding classical system. For instance, in constructron theory, Marletto et al. define a classical system based on the concept of cloning [38, 17], while Krisnand et al. define it in relation to commutativity [39]. Various definitions of classical systems are interconnected and may exhibit varying degrees of effectiveness in different analytical contexts.
In this paper, by “classical” we mean precisely a system that has only one single observable (single basis). The other observables of a classical system are functions of this observable [36]. The definition of a classical system here does not only require that its corresponding physical states are all perfectly distinguishable, but it also implies that it is restricted from interacting with any matter, i.e., only these commutative observables (one basis) are involved in the interaction [19, 40].
A natural consequence of the definition of classical systems is that classical systems are only classically correlated with any system (including quantum systems). Since a system without two or more than two non-commutative observables is not able to establish a quantum correlation with a quantum system, such as a quantum discord and quantum entanglement [41]. Besides, one may infer that a classical system can perform different control operations on quantum systems, while quantum systems cannot. Because if different orthogonal states of the quantum system control different operations of the classical system, then the superposition of these orthogonal states will lead to the transmission of quantum information to the classical system’s thereby creating quantum correlations, which contradicts the definition of classicality. Based on these properties, we will show how to construct a general dynamics of a hybrid system in an operationally defined way.
II Quantum-classical Dynamics
Classical correlations manifest in the majority of physical dynamic processes. These correlations can arise between quantum-quantum systems, classical-classical systems, or quantum-classical systems. In the subsequent discussion, our attention will be directed towards exploring the quantum-classical hybrid systems. Before proceeding, we will provide the necessary notations and definitions for both quantum and classical systems.
For a quantum system, we utilize the density matrix in Hilbert space to represent a quantum state. On the other hand, for a classical system, we do not make any assumptions about a specific model. To maintain generality, we denote a pure classical state as “$\text{State}_{i}$”, where each unique value of $i$ represents a perfectly distinguishable state. Naturally, for a mixed classical state, it can be defined as follows: $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$, where $p_{i}$ is the probability of different pure states “$\text{State}_{i}$”.
As classical systems cannot be quantum correlated with quantum systems, their composite states exhibit solely classical correlations. Suppose the initial state of the hybrid system lacks classical correlation, following a quantum-classical interaction, the resulting state generally evolves and can be expressed in the following general form:
$\wp_{\mathcal{C}}\oplus\varrho\rightarrow\sum_{i}q_{i}\wp_{{\mathcal{C}}i}\oplus\varrho_{i},$ (1)
where $\{q_{i}\}$ is the probability distribution of classical correlation and $\wp_{{\mathcal{C}}i}=\sum_{j}p_{j}\text{State}_{i}$. Historically, the evolutionary state of a hybrid system is called the hybrid density operator [42, 43], which satisfying an overall normalization condition and preserving the form of the state [48].
An operational definition of quantum-classical dynamics $\Theta_{\mathcal{QC}}$ can be derived based on the limitations of classical correlations and the properties of classical systems. As depicted in Figure 1, we provide a general operational definition for quantum-classical dynamics over infinitesimal time intervals. In this context, we refrain from imposing any specific theoretical framework on quantum-classical dynamics and instead focus on its operational decomposition, emphasizing how it generates classical correlations. It’s important to note that while we are examining the infinitesimal time evolution of hybrid systems, this approach possesses the generality for straightforward extension to continuous time dynamics (see Supplementary Material).
Figure 1: Operational definition of quantum-classical dynamics in infinitesimal time.
Without loss generality, we suppose the initial state of the hybrid system is uncorrelated, denoted as $\chi_{\mathcal{QC}}=\wp_{\mathcal{C}}\oplus\varrho$, where $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$. Generally, one can divide the quantum-classical dynamics into two distinct parts. Firstly, there is the map $V$, which operates on the classical state $\wp_{\mathcal{C}}$. Consequently, $V$ transforms the classical state $\wp_{\mathcal{C}}$ into $V\cdot\wp_{\mathcal{C}}=V\cdot\sum_{i}p_{i}\text{State}_{i}=\sum_{i}q_{i}\text{State}_{i}$ and the hybrid state now becomes $\sum_{i}q_{i}\text{State}_{i}\oplus\varrho$. The second part involves the establishment of a control channel $\Lambda_{i}$ from the classical sector to the quantum sector, from which the classical correlations of the states of the composite system originate. The specific process of the second part can be expressed as:
$\sum_{i}q_{i}\text{State}_{i}\oplus\varrho\rightarrow\sum_{i}q_{i}\text{State}_{i}\oplus\varrho_{i},$ (2)
where $\rho_{i}=\Lambda_{i}\cdot\varrho$. Therefore, one has $\Theta_{\mathcal{QC}}\cdot\chi_{\mathcal{QC}}=\chi_{\mathcal{QC}}^{\prime}=\sum_{i}q_{i}\text{State}_{i}\oplus\varrho_{i}$. Note that (a) any local operation on a quantum system can be incorporated into the $\Lambda$; (b) the dynamics $\Theta_{\mathcal{QC}}$, $V$ and $\Lambda$ don’t need to be (input) state-independent local operations and control operations, but rather state-dependent. Since for a general hybrid dynamics $\Theta_{\mathcal{QC}}$ one do not exclude the possibility that it is a nonlinear mapping (no linearity and reversibility need to be assumed). It doesn’t matter what the nature of the hybrid dynamics is, as long as it ends up embodying the state of classical correlation. Thus, in general, $\Theta_{\mathcal{QC}}=\Theta_{\mathcal{QC}}(\chi_{\mathcal{QC}}),V=V(\chi_{\mathcal{QC}})$, $\Lambda_{i}=\Lambda(\chi_{\mathcal{QC}})$ and $q_{i}=q_{(}\chi_{\mathcal{QC}})$, which all may be the function of $\chi_{\mathcal{QC}}$. For convenience, we abbreviate them as $V$, $\Lambda_{i}$ and $q_{i}$ below.
The above operational definition of hybrid system requires that a legal physical state is mapped to another physical state at any moment [42, 43]. An alternative approach to comprehend this decomposition is to liken the classical state $\{\text{State}_{i}\}$ to a set of orthogonal bases $\{|i\rangle\langle i|\}$. This analogy arises from the fact that the classical system only possesses one basis (one context), and any arbitrary mapping of its interactions with the quantum system can always be broken down into local operations and control operations of the classical to the quantum system since arbitrary two-body operations can always be decomposed using local operations and control operations [49]. As we mentioned before, instead of assuming linearity, we assume that the control of a classical system by a quantum system is not allowed, hence the most general form of quantum -classical dynamics over infinitesimal time intervals must be one that cascades from the circuit shown in Figure 1 (see Supplementary Material for more details of continuous dynamics).
III Observables in Hybrid systems
Different physical theories, such as classical mechanics and quantum mechanics, may exhibit incompatible mathematical structures. However, they share a common feature: the ability to describe physical systems using expectation values. In terms of expectation values, both classical and quantum theories yield real numbers. The validity of a theory is determined by whether its predicted expectations match those observed experimentally. Therefore, expectation values of observables (dynamic variables) serve as a bridge connecting different theories. From a measurable perspective, a general description of expectation value of observable is required.
The attributes of a physical system, specifically the expectation values of observables, are jointly determined by the state and observables. Now the relationship between the state of a physical system and it’s observable can be succinctly described as
[…]
IV A No go theorem
Here we present a no-go theorem that demonstrates the impossibility of any quantum-classical dynamics causing a change in the local observable of either a classical or quantum system while preserving the conservation of the global observable. In other words, the conservation law imposes limitations on the exchange of corresponding local observables between classical and quantum systems. The no-go theorem is as follows:
Theorem 1
Any quantum classical dynamics that satisfies the decomposition in figure 1 can not change the local observable $\langle O_{\mathcal{C/Q}}\rangle$ of either a classical or quantum system under the conser vation law of a global quantity of the hybrid system $\mathcal{C}\oplus\mathcal{Q}$. ${}_{\Box}$
Let’s see how this theorem is derived. Suppose there is a conserved observable $O=O_{\mathcal{C}}+O_{\mathcal{Q}}$ of a quantum-classical hybrid system, satisfying
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t+\Delta t},$ (4)
where $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle=\{\chi_{\mathcal{QC}},O_{\mathcal{C}}\}+\{\chi_{\mathcal{QC}},O_{\mathcal{Q}}\}=\langle O_{\mathcal{C}}\rangle+\langle O_{\mathcal{Q}}\rangle$. The conservation law dictates that equation (4) must hold at all times. This implies that, as illustrated in Figure 1, every moment of the dynamical evolution of the hybrid system must adhere to the conservation law. In other words, both the preceding local operation and the subsequent control operation must independently comply with the conservation of global observable $O$.
Without loss generality, suppose the initial state of hybrid system has no classical correlation, that is $\chi_{\mathcal{QC}}=\wp_{\mathcal{C}}\oplus\varrho$, where $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$. Immediately, one may calculate the corresponding observable $O$ as
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{0}}=\sum_{i}p_{i}\{\text{State}_{i},O_{\mathcal{C}}\}+\text{Tr}(\varrho O_{\mathcal{Q}}).$ (5)
As we have analyzed before, after the operation $V$, the hybrid state become $V\cdot\varrho\oplus\wp_{\mathcal{C}}=\sum_{i}q_{i}\text{State}_{i}\oplus\varrho$. We have
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{1}}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}+\text{Tr}(\varrho O_{\mathcal{Q}}).$ (6)
Similarly, the control operation $\Lambda_{i}$ in Figure 1 changes the state from $\sum_{i}q_{i}\text{State}_{i}\oplus\varrho$ to $\sum_{i}q_{i}\text{State}_{i}\oplus\varrho_{i}$ with
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2}}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}+\sum_{i}q_{i}\text{Tr}(\varrho_{i}O_{\mathcal{Q}}),$ (7)
where $\rho_{i}=\Lambda_{i}\cdot\varrho$. Since $O$ is conserved at any time, there must be $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{0}}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{1}}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2}}$. Obviously, we have $\sum_{i}p_{i}\{\text{State}_{i},O_{\mathcal{C}}\}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}$ and
$\langle O_{\mathcal{Q}}\rangle_{t_{0}}=\text{Tr}(\varrho O_{\mathcal{Q}})=\sum_{i}q_{i}\text{Tr}(\varrho_{i}O_{\mathcal{Q}})=\langle O_{\mathcal{Q}}\rangle_{t_{2}}.$ (8)
It is not difficult to find that the evolution from $t_{0}$ to $t_{n}$ still has $\langle O_{\mathcal{Q}}\rangle_{t_{0}}=\langle O_{\mathcal{Q}}\rangle_{t_{n}}$ (see Supplementary materials). This states that a hybrid dynamical model of the quantum-classical system cannot make the local $\langle O_{\mathcal{Q/C}}\rangle$ change while satisfying the conservation of global $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle$.
We will now show that the interaction of two quantum systems is exchangeable up to local observables in the case of conservation laws, which is fundamentally different from the hybrid dynamics. In quantum theory, the conserved observable $O$ satisfies
$\langle[O_{\mathcal{Q}_{1}}+O_{\mathcal{Q}_{2}},U_{\mathcal{Q}_{1}\mathcal{Q}_{2}}]\rangle=0,$ (9)
where $U_{\mathcal{Q}_{1}\mathcal{Q}_{2}}$ is the unitary evolution of composite system $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$. In general, $\langle[O_{\mathcal{Q}_{i}},U_{\mathcal{Q}_{1}\mathcal{Q}_{2}}]\rangle\neq 0$, where $i=1,2$. That is the local observable $O_{\mathcal{Q}_{i}}$ is generally not a conserved quantity. Therefore, one may have
$\langle O_{\mathcal{Q}}\rangle_{t}\neq\langle O_{\mathcal{Q}}\rangle_{t+\Delta t}.$ (10)
Based on the inconsistency between the predictions of quantum-classical and full quantum dynamics, one immediately has
Corollary 1
If the local observable $\langle O_{S}\rangle$ of a quantum system $S$ is changed after interacting with an unknown matter $E$, under the conservation law of a global quantity $\langle O_{S}\rangle+\langle O_{E}\rangle$ of the global system $SE$ and assuming that the dynamics satisfies the decomposition in figure 1, then $E$ must be non-classical. ${}_{\Box}$
That is, subject to a specific conservation law, if a matter of unknown nature can induce a change in the corresponding local observable of a quantum system, it must imply that this unknown matter must be quantumlike. Therefore, one may witness the nature of an unknown matter even though the matter are inaccessible.
V Quantumness of gravitational field
Figure 2: Momentum change of two quantum matters in the interaction of gravitational fields. (a) At the moment of $T=t$, the initial positions of the two stationary masses are $x_{S}$ and $x_{M}$, and the momentum are $\langle P_{\mathcal{Q/C}}\rangle=p_{S}=p_{M}=0$. (b) At the moment of $T=t+\Delta t$, the positions of masses $S$ and $M$ have changed, and their respective momentum changes in the same amount and in opposite directions while the whole satisfies the conservation of momentum.
In this section, we will argue the quantumlike nature of gravity based on the no-go theorem we proposed. Before that, we introduce the assumption of locality. Locality is defined here as any interaction between objects at a distance that is mediated through a field, rather than being instantaneous. This asserts that the gravitational interaction between objects $A$ and $B$ is mediated through a gravitational field. That is there is no direct interaction between objects $A$ and $B$.
Now we denote the gravitational field labeled as $E$, which is primarily generated by a massive object denoted as $M$. As depicted in Figure 2 (a), at $T=t_{0}$, we suppose there are two stationary quantum masses labeled as $S$ and $M$ respectively. Based on our existing observations of the universe, these masses may mutually attract to each other as a consequence of the gravitational field $E$ mediating their interaction. While the total momentum of the system remains conserved at zero, the individual momentum of system $S$ and system $M$ increases in opposite directions, as shown in Figure 2 (b).
Let us analyze how the predictions of the classical gravitational field contradict the above description of the phenomenon. Here we delve into the discussion of the gravitational field in two cases. In the first case, the gravitational field $E$ is considered to be classical. Due to the local nature of interaction, the interaction survives only in $ES$ and $EM$. Based on our no-go theorem, and assuming momentum conservation, the gravitational field of classical nature $E$ cannot induce any change in the momentum $P$ of quantum matter $S$. In other words,
$\langle P_{S}\rangle_{t}=\langle P_{S}\rangle_{t+\Delta t},$ (11)
where $\langle P_{S}\rangle_{t}=\text{Tr}(P\varrho)$ and $\langle P_{S}\rangle_{t}=\text{Tr}(P\varrho^{\prime})$. Moving on to the second case, let’s now consider $E$ is a quantum gravitational field. In this scenario, the Hamiltonian of composed system can be expressed as $H_{SME}=H_{S}+H_{E}+H_{M}+H_{SE}+H_{EM}$. Remarkably, the Hamiltonian allows us to meet the momentum conservation criteria, ensuring that $\langle[U_{SEM},P_{S}+P_{E}+P_{M}]\rangle=0$ while $\langle[U_{SEM},P_{S/E/M}]\rangle\neq 0$, where $U_{SEM}=\text{exp}(-iH_{SEM}T/\hbar)$. That is, one may has
$\langle P_{S}\rangle_{t}\neq\langle P_{S}\rangle_{t+\Delta t}.$ (12)
Compare to the classical gravitational field, the quantum one has ability to induce changes in the momentum of quantum system for both $S$ and $M$. Consequently, if a particle is observed to change momentum after being subjected to a gravitational field, then the gravitational field must be quantum.
Let now suppose $M$ to be a huge massive object, i.e. Earth or a metal sphere, representing quantum matter that adheres to the classical limit, as shown in Figure 2. According to our argument, under the principle of momentum conservation, any momentum change in the quantum mass $S$ resulting from the influence of $M$ mediated through the gravitational field is viewed as a sufficient evidence for the quantum nature of gravity. Actually, there are lots of experimental observation to support our argument. For example, both the free fall of a quantum particle and the displacement of particle caused by gravitational interaction of the metal sphere in the recent gravitational Aharonov-Bohm effect [50] offer compelling indications of the quantumness of gravity.
VI Discussions
While we have argued for the non-classical nature of gravity in terms of conservation of momentum, a natural question is the why conservation of momentum is special? Is it possible to derive the same conclusion using other conservation laws, such as conservation of energy. The answer is yes. But before doing that, here we need to revisit the definition of energy. When we consider the gravitational field as an interacting object, the energy of the particle may be just its own free energy and may not include the gravitational potential energy. Because the potential energy of a particle comes from interactions (just ignore other large systems and fields), this is an “approximate” definition from classical physics. This may applies to classical mechanics and semi-classical quantum theory, but not to full quantum theory.
Let’s explore the concept further: Imagine a particle’s energy, defined as free energy (kinetic energy). Our above analysis asserts that if the kinetic energy of a particle increases following an interaction with a gravitational field, this significant observation implies that gravity behaves in a manner inconsistent with classical objects.
Evidently, the free-fall of a particle provides compelling evidence for the quantum nature of gravity, assuming the principle of energy conservation holds true (similar to the argument of conservation of momentum, except that the conserved quantity is replaced by energy in the argument). Furthermore, when we scrutinize other conserved quantities, if a change consistent with the observed quantity occurs in a probing system, it logically follows that the entity responsible for mediating this interaction, such as the gravitational field, must adhere to the principles of quantum mechanics. Obviously, with respect to quantum-classical hybrid dynamics based on Hamiltonian, this conclusion also holds [51].
VII conclusions
We have developed a general framework for quantum-classical hybrid dynamics, offering an approach that is model-independent. This framework is predicated on the fundamental premise that any classical system can only maintain classical correlations with a quantum system.
Within this framework, we present a no-go theorem that pertains to systems subject to a conservation law. The no-go theorem explores a fundamental limitation of classical system to interact with a quantum system, assuming that the interaction satisfies the decomposition property set out in figure 1, i. e. any classical system can not induce a change of the local observable of quantum system while a global conserved quantity of hybrid system is assumed. Based on this no-go theroem, our exploration delves into the quantum nature of gravity, unveiling a profound connection that exists between momentum conservation and the inherent quantum aspects of gravity. It is worth noting that a similar argument may be extended to the conservation of energy.
Additionally, our framework underscores the pivotal role played by expectation values in facilitating a holistic comprehension of the physical universe. This transcends the limitations imposed by disparate mathematical structures and serves as a bridge, connecting various physical theories into a unified description.
While our argument regarding the quantum nature of gravity remains within a non-relativistic context, its generality and validity persist. We expect to discover more relationships between fundamental principles and gravity, leading to a deeper understanding of the nature of spacetime.
Acknowledgements.
T. F. thanks the support of QICI of the University of Hong Kong, the National Natural Science Foundation of China, and is grateful to Tian Zhang for helpful discussions. C. M. and V.V. thank the Gordon and Betty Moore Foundation. This publication was also made possible through the support of the ID 61466 grant from the John Templeton Foundation, as part of the The Quantum Information Structure of Spacetime (QISS) Project (qiss.fr). The opinions expressed in this publication are those of the Authors and do not necessarily reflect the views of the John Templeton Foundation.