### v3

# Conservation Laws and the Quantization of Gravity

Abstract
Adopting general frameworks for quantum-classical dynamics, we analyze the interaction between quantum matter and a classical gravitational field. We point out that, assuming conservation of momentum or energy, and assuming that the dynamics obeys Hamiltonian formalism or a particular decomposition property set out in the paper, the classical gravitational field cannot change the momentum or energy of the quantum system, whereas the quantum gravitational field can do so. Drawing upon the fundamental relationship between conservation laws and the quantum properties of objects, our analysis offers new perspectives for the study of quantum gravity and provides a novel interpretation of existing experimental observations, such as free fall.
The quantization of gravity, a fundamental issue of modern theoretical physics, has captivated the imaginations of physicists for over a century. This profound quest seeks to reconcile two pillars of contemporary physics: general relativity and quantum theory. Yet, it is hard to unify them into a unique corpus of laws. Several approaches, such as field theory, “geometrization” of quantum mechanics and string theory, have emerged in the pursuit of reconciling quantum theory with general relativity [1, 2, 3]. However, all these approaches face significant theoretical and conceptual challenges [4, 5], which have prompted some to propose the fundamental classical nature of gravity [6]. Further, some have even argued that quantizing gravity is not needed after all [7] or that gravity may not even be a fundamental force [8, 9]. Whether or not the quantisation of gravity is necessary has become highly contentious [10, 11, 12, 13, 15, 14], primarily due to the enduring difficulty in acquiring direct supporting evidence [16, 17].
As observed by Feynman during the 1950 Chapel Hill conference, it was pointed out that the physical state of the composite system exhibits distinct properties depending on whether the mass is entangled with gravity or merely classically correlated with it [20]. This highlights the importance of the superposition of large masses in the study of quantum gravity. Recently, there has been a revival of Feynman’s idea , which highlights gravitationally induced entanglement between two massive particles provide a strong evidence on whether the gravitational field is mediated by the transfer of quantum information [18, 19]. This is clearly sufficient evidence for quantum effects in gravity. Unfortunately, gravitationally induced entanglement is extremely weak, and even though there are some advanced theoretical analyses [21, 22, 23, 24] and experimental protocols [25, 26, 27, 28, 29, 30] available, with the current experimental techniques and advances, it is not yet time to observe this effect.
The progress outlined above based on quantum information perspective is truly enlightening, even though the pursuit of quantum entanglement induced by gravity remains a formidable challenge. A natural question arises: Are the existing experimental observations of gravity sufficient to analyze the quantum nature of gravity? In this paper, we unequivocally affirm this possibility.
In historical context, Eppley and Hannah endeavored to construct a logical argument challenging the classical model of gravity’s compatibility with the conservation of momentum [10]. It is essential to acknowledge that their work predominantly consisted of a thought experiment and lacked empirical observations. Subsequently, their assertions have faced widespread skepticism and have been refuted by subsequent research [11, 13]. Moreover, other gedankenexperiments have been conducted, highlighting the perplexing paradoxes of causality and complementarity arising from classical gravity [31, 32]. Indeed, the introduction of some basic principles and physical laws helps us to analyse the non-classical nature of matter, even in the absence of entanglement witnesses [33, 34, 35, 36].
Inspired by [34, 35], our work aims to delve into the intrinsic relationship between conservation laws and the underlying nature of gravity. We first investigate general quantum-classical dynamics based on Hamiltonian formalism, pointing out any quantum-classical hybrid dynamics based on the Hamiltonian formalism cannot have backreaction from the quantum sector to the classical sector. This indicates all observables of classical systems are conserved quantities in this model. For the case beyond Hamiltonian formalism, we shall aim at relaxing some of the assumptions in [34], e.g. linearity of the overall dynamics, by adding another constraint on the dynamics, outlined in section II. To accomplish this, we harness information-theoretic methodologies to construct a general framework of hybird systems that amalgamates quantum and classical dynamics into a unified whole [37]. Within these two frameworks (Hamiltonian formalism and the latter one), we present no-go theorems, establishing a fundamental limitation. They stipulate that any quantum-classical dynamics cannot induce alterations in the local observables of either classical or quantum systems, all the while maintaining the conservation of global observables. Our results bring to light a profound revelation: when considering the conservation of either momentum or energy, the classical gravitational field is unable to effect changes in the momentum or energy of a quantum system. In other words, in these frameworks, conservation laws are sufficient but not necessary conditions for the quantization of gravity. This proposition stands in contrast to established experimental observations, such as the free fall experiment. Conversely, our arguments compellingly demonstrates that the quantum gravitational field possesses the capability to enact such changes, offering compelling evidence for the quantum nature of gravity.
I Definition of a Classical system
A fundamental distinction between classical and quantum systems lies in their contextual nature. In this paper, by “classical” we mean precisely a system that has only one single observable (single basis, e.g. $\{|i\rangle\langle i|\}$). The other observables of a classical system are functions of this observable [36, 14]. The definition of a classical system here does not only require that its corresponding physical states are all perfectly distinguishable, but it also implies that it is restricted from interacting with any matter, i.e., only these commutative observables are involved in the interaction [39, 19, 40].
A natural consequence of the definition of classical systems is that classical systems are only classically correlated with any system, including quantum systems [41]. Since a system without two or more than two non-commutative observables is not able to establish a quantum correlation with a quantum system. Besides, one may infer that a classical system can perform different control operations on quantum systems, while quantum systems cannot [42].
II Hamiltonian formalism of a genuine quantum-classical hybrid system
Historically, extensive discussions have revolved around hybrid quantum-classical systems [43, 47, 45], often employing models rooted in the Hamiltonian framework proposed by Aleksandrov [43] and Koopman [45]. Aleksandrov’s framework primarily serves to analyze the manner in which classical and quantum systems revert to their respective equations of motion following their interaction. In contrast, Koopman’s framework introduces the superselection rule, primarily utilized to address the measurement problem when coupling classical devices with quantum matter [46]. Nonetheless, it’s important to note that these frameworks for hybrid systems inherently maintain a quantum nature. While the classical-like behavior is assured within the evolutionary state, the entire dynamics of hybrid systems necessitate the introduction of noncommutative operators in the classical component of the system, including its state, involved in the coupling.
As articulated by Diósi, the establishment of a consistent hybrid dynamics necessitates a process termed “requantization” for the classical system [48, 44]. In the context of such hybrid models, these classical systems are appropriately referred to as quasi-classical systems, which fundamentally quantum [48]. Similarly, a prominent framework of quantum-classical interaction proposed by Hall and Reginatto’s also based on Hamiltonian formalism [49]. Most of previous discussions on hybrid system models, encompassing Hamiltonian formalism, mainly sought a dynamical equation that can be regressed to their respective scales. These discussions, however, are not suited for testing whether gravity necessitates quantization. It’s noteworthy to mention Oppenheim’s recent work on a post-quantum theory of classical gravity [53, 54], which builds upon an extension of Aleksandrov’s methodology. It may be seen as an effective quantum theory of “classical” gravity based on the theory of open quantum systems. Nonetheless, as we have previously discussed, it should be distinguished from a genuine quantum theory of classical gravity, as it incorporates noncommutative operators into the classical system during the construction of the global system Hamiltonian.
Here we give a simple analysis for the most general form of coupling between classical systems (e.g. gravitational fields) and quantum matter in the Hamiltonian framework based on the strict definition of classical systems. Without loss generality, we suppose the unique basis of classical system is $\{|i\rangle\langle i|\}$. Generally, the Hamiltonian of quantum-classical hybrid system is
$\mathbf{H}=H_{\mathcal{C}}+H_{\mathcal{Q}}+\sum_{i}\gamma_{i}|i\rangle\langle i|\otimes S_{i},$ (1)
where $H_{\mathcal{C}}$ ($H_{\mathcal{C}}=\sum_{i}\epsilon_{i}|i\rangle\langle i|$) and $H_{\mathcal{Q}}$ are the local Hamiltonian of the classical system and quantum system respectively. The last term $\sum_{i}\gamma_{i}|i\rangle\langle i|\otimes S_{i}$ is the interaction Hamiltonian of classical system and quantum system, where $S_{i}$ is the Hermitian operator of the quantum system. Obviously, any observable of the classical system in this model is an invariant (no backaction from quantum sector to classical one) since
$[O_{\mathcal{C}},\mathbf{H}]=0$ (2)
where $O_{\mathcal{C}}=\sum_{i}o_{i}|i\rangle\langle i|$. Immediately, this leads a simple no-go theorem:
Theorem 1
Any quantum-classical hybrid dynamics based on the Hamiltonian formalism cannot have backreaction from the quantum sector to the classical sector.
This theorem indicates all observables of classical systems are conserved quantities in the hybrid quantum-classical model with Hamiltonian formalism.
This result is consistent with the classical limits of quantum physics. As in the Markovian open-system environment, a quantum particle’s reaction to its environment is so negligible that it cannot change the state of the environment (such as thermal baths and the Earth’s gravitational field). In this case, all observables of the environment can be regarded as invariant. We note that this conclusion may compatible with the recent work of Galley et al. [55]. There is one possible exception to this, regarding the measurement of quantum particles by detectors. This measurement process it has to be considered in terms of a full quantum model, such that photons are absorbed by the detector and eventually amplified and converted into electrical signals to be observed. This process obviously requires that the quantum particles have a reaction on the measurement device such that it changes the state of the measurement device. Thus the problem of quantum measurement must be a quantum issue and not one that can be explained by the quantum-classical model.
On the other hand, the classical reaction against the quantum is obvious. Generally, one has
$[O_{\mathcal{Q}},\mathbf{H}]\neq 0.$ (3)
As an example, the Colella-Overhauser-Werner (COW) experiment illustrates how the Earth’s gravitational field (quasiclassical system) induces quantum interference of neutrons (quantum system) [56, 57]. For different paths $|x_{1}\rangle$ and $|x_{2}\rangle$ of the neutron interferometer, the neutron coherently couples to the gravitational field at different locations (corresponding to couple different classical state $|i\rangle\langle i|$), producing different phases (depending on $\gamma_{i}$). As is well known, this gravitationally induced phase is dependent on the difference in potential energy between the two paths, so $\gamma_{i}$ is a parameter on the potential energy. Because the COW experiment only considers gravitationally induced phases and does not take into account the exchange of momentum or energy between the gravitational field and the quantum particles, one cannot rule out the classical model of gravity purely from the COW experiment.
The argument for classical gravity is untenable when one deals with experiments such as free fall. Indeed, if we recognise conservation laws such as the conservation of energy and momentum, then the fact that an object (a macroscopic quantum system) acquires energy or momentum from a gravitational field must preclude the possibility of gravity being classical. The gravitational field that mediates this exchange of energy or momentum must possess two and more non-commutative dynamics variables (as we show in the following). From a testability perspective, one may summarize the following corollary:
Corollary 1
If the local observable $\langle O_{S}\rangle$ of a quantum system $S$ is changed after interacting with an unknown matter $E$, under the conservation law of a global quantity $\langle O_{S}\rangle+\langle O_{E}\rangle$ of the global system $SE$ and assuming that the dynamics satisfies the Hamiltonian formalism, then $E$ must be non-classical.
Note that the above conclusion applies not only to the case where the global evolution of quantum-classical systems is $e^{-i\mathbf{H}t}$, but also to the mapping $f(\mathbf{H})$ of any function of $\mathbf{H}$, since
$[O_{\mathcal{C}},f(\mathbf{H})]=0.$ (4)
So for the hybrid dynamics with Hamiltonian formalism, the gravitational field must also be quantum, otherwise there is no backreaction of the quantum system on the classical system or it violates the conservation laws.
III Quantum-classical Dynamics Beyond Hamiltonian formalism
Classical correlations manifest in the majority of physical dynamic processes. These correlations can arise between quantum-quantum systems, classical-classical systems, or quantum-classical systems. In the subsequent discussion, our attention will be directed towards exploring general quantum-classical hybrid systems beyond Hamiltonian formalism. Before proceeding, we will provide the necessary notations and definitions for both quantum and classical systems.
[…]
As classical systems cannot be quantum correlated with quantum systems, their composite states exhibit solely classical correlations. Suppose the initial state of the hybrid system lacks classical correlation, following a quantum-classical interaction, the resulting state generally evolves and can be expressed in the following general form:
$\wp_{\mathcal{C}}\oplus\varrho\rightarrow\sum_{i}q_{i}\wp_{{\mathcal{C}}i}\oplus\varrho_{i},$ (5)
where $\{q_{i}\}$ is the probability distribution of classical correlation and $\wp_{{\mathcal{C}}i}=\sum_{j}p_{j}\text{State}_{i}$. Historically, the evolutionary state of a hybrid system is called the hybrid density operator [43, 44], which satisfying an overall normalization condition and preserving the form of the state.
An operational definition of quantum-classical dynamics $\Theta_{\mathcal{QC}}$ can be derived based on the limitations of classical correlations and the properties of classical systems. As depicted in figure 1, we provide a general operational definition for quantum-classical dynamics over infinitesimal time intervals. In this context, we refrain from imposing any specific theoretical framework on quantum-classical dynamics and instead focus on its operational decomposition, emphasizing how it generates classical correlations.
Figure 1: Operational definition of a class of quantum-classical dynamics. (a) Quantum-classical dynamics of multi-time evolution. Each time interval is infinitesimal, that is, $t_{i}-t_{i+1}=\delta t$. The dynamics of a hybrid system consists of sequentially infinitesimal slices of evolution; (b) The quantum-classical dynamic in a single interval (infinitesimal time).
Without loss generality, here we analyze the hybrid dynamics of an infinitesimal time slice, as in figure 1 (b), while the proof of continuous time is put into appendixes. We suppose the initial state of the hybrid system is uncorrelated, denoted as $\chi_{\mathcal{QC}}=\wp_{\mathcal{C}}\oplus\varrho$, where $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$. Generally, as shown in figure 1 (b), one may divide the quantum-classical dynamics into two distinct parts. Firstly, there is a global map which is effectively equivalent to the local action $V$ operating on the classical state $\wp_{\mathcal{C}}$. Consequently, $V$ transforms the classical state $\wp_{\mathcal{C}}$ into $V\cdot\wp_{\mathcal{C}}=V\cdot\sum_{i}p_{i}\text{State}_{i}=\sum_{i}q_{i}\text{State}_{i}$ and the hybrid state now becomes $\sum_{i}q_{i}\text{State}_{i}\oplus\varrho$. The second part involves the establishment of a control channel $\Lambda_{i}$ from the classical sector to the quantum sector, from which the classical correlations of the states of the composite system originate. The specific process of the second part can be expressed as:
$\sum_{i}q_{i}\text{State}_{i}\oplus\varrho\rightarrow\sum_{i}q_{i}\text{State}_{i}\oplus\varrho_{i},$ (6)
where $\rho_{i}=\Lambda_{i}\cdot\varrho$. Therefore, one has $\Theta_{\mathcal{QC}}\cdot\chi_{\mathcal{QC}}=\chi_{\mathcal{QC}}^{\prime}=\sum_{i}q_{i}\text{State}_{i}\oplus\varrho_{i}$. Note that (a) any local operation on a quantum system can be incorporated into the $\Lambda$; (b) the dynamics $\Theta_{\mathcal{QC}}$, $V$ and $\Lambda$ don’t need to be (input) state-independent local operations and control operations, but rather state-dependent. Since for a general hybrid dynamics $\Theta_{\mathcal{QC}}$ one do not exclude the possibility that it is a nonlinear mapping (no linearity and reversibility need to be assumed). It doesn’t matter what the nature of the hybrid dynamics is, as long as it ends up embodying the state of classical correlation. Thus, in general, $\Theta_{\mathcal{QC}}=\Theta_{\mathcal{QC}}(\chi_{\mathcal{QC}}),V=V(\chi_{\mathcal{QC}})$, $\Lambda_{i}=\Lambda(\chi_{\mathcal{QC}})$ and $q_{i}=q_{(}\chi_{\mathcal{QC}})$, which all may be the function of $\chi_{\mathcal{QC}}$. For convenience, we abbreviate them as $V$, $\Lambda_{i}$ and $q_{i}$ below.
The above operational definition of hybrid system requires that a legal physical state is mapped to another physical state at any moment [43, 44]. An alternative approach to comprehend this decomposition is to liken the classical state $\{\text{State}_{i}\}$ to a set of orthogonal bases $\{|i\rangle\langle i|\}$. This analogy arises from the fact that the classical system only possesses one basis, and any arbitrary mapping of its interactions with the quantum system can always be broken down into local operations and control operations of the classical to the quantum system [50]. As we mentioned before, instead of assuming linearity, we assume that the control of a classical system by a quantum system is not allowed, hence the most general form of quantum-classical dynamics over infinitesimal time intervals must be one that cascades from the circuit shown in Figure 1.
III.1 Observables in Hybrid systems
The attributes of a physical system, specifically the expectation values of observables, are jointly determined by the state and observables. Here the relationship between the state of a physical system and it’s observable can be succinctly described as
[…]
In a classical theory, the expectation value of a classical system can be described as $\{\wp_{\mathcal{C}},O\}\equiv\sum_{i}p_{i}\{\text{State}_{i},O\}$, where $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$. In particular, if the $\text{State}_{i}$ is the eigenstate of eigenvalue $O_{i}$, then $\sum_{i}p_{i}\{\text{State}_{i},O\}=\sum_{i}p_{i}O_{i}$, which is similar to the definition of legitimate expectation value $\langle O\rangle$. For a hybrid quantum-classical system, the expectation value of observable is naturally defined as follows $\langle O\rangle=\{\chi_{\mathcal{QC}},O\},$ where $\chi_{\mathcal{QC}}$ is the quantum-classical hybrid state defined above. If $O$ is an additive observable, i. e. $O=O_{\mathcal{C}}+O_{\mathcal{Q}}$, $O_{\mathcal{C}}+O_{\mathcal{Q}}=\{\chi_{\mathcal{QC}},O_{\mathcal{C}}\}+\{\chi_{\mathcal{QC}},O_{\mathcal{Q}}\}$ ( See Appendixes for more more details of observables in hybrid systems).
III.2 A No go theorem
Here we present a no-go theorem that demonstrates the impossibility of a class of quantum-classical dynamics (satisfying the decomposition in figure 1) causing a change in the local observable of either a classical or quantum system while preserving the conservation of the global observable. In other words, the conservation law imposes limitations on the exchange of corresponding local observables between classical and quantum systems. The no-go theorem is as follows:
Theorem 2
Any quantum classical dynamics that satisfies the decomposition in figure 1 can not change the local observable $\langle O_{\mathcal{C/Q}}\rangle$ of either a classical or quantum system under the conservation law of a global quantity of the hybrid system $\mathcal{C}\oplus\mathcal{Q}$.
Let’s see how this theorem is derived. Suppose there is a conserved observable $O=O_{\mathcal{C}}+O_{\mathcal{Q}}$ of a quantum-classical hybrid system, satisfying
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t+\Delta t},$ (8)
where $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle=\{\chi_{\mathcal{QC}},O_{\mathcal{C}}\}+\{\chi_{\mathcal{QC}},O_{\mathcal{Q}}\}=\langle O_{\mathcal{C}}\rangle+\langle O_{\mathcal{Q}}\rangle$. The conservation law dictates that equation (8) must hold at all times. This implies that, as illustrated in Figure 1, every moment of the dynamical evolution of the hybrid system must adhere to the conservation law.
Without loss generality, suppose the initial state of hybrid system has no classical correlation, that is $\chi_{\mathcal{QC}}=\wp_{\mathcal{C}}\oplus\varrho$, where $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$. Immediately, one may calculate the corresponding observable $O$ as
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{0}}=\sum_{i}p_{i}\{\text{State}_{i},O_{\mathcal{C}}\}+\text{Tr}(\varrho O_{\mathcal{Q}}).$ (9)
As we have analyzed before, after the operation $V$, the hybrid state become $V\cdot\varrho\oplus\wp_{\mathcal{C}}=\sum_{i}q_{i}\text{State}_{i}\oplus\varrho$. We have
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{1}}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}+\text{Tr}(\varrho O_{\mathcal{Q}}).$ (10)
Similarly, the control operation $\Lambda_{i}$ in Figure 1 (b) changes the state from $\sum_{i}q_{i}\text{State}_{i}\oplus\varrho$ to $\sum_{i}q_{i}\text{State}_{i}\oplus\varrho_{i}$ with
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2}}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}+\sum_{i}q_{i}\text{Tr}(\varrho_{i}O_{\mathcal{Q}}),$ (11)
where $\rho_{i}=\Lambda_{i}\cdot\varrho$. Since $O$ is conserved at any time, there must be $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{0}}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{1}}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2}}$. Obviously, we have $\sum_{i}p_{i}\{\text{State}_{i},O_{\mathcal{C}}\}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}$ and
$\langle O_{\mathcal{Q}}\rangle_{t_{0}}=\text{Tr}(\varrho O_{\mathcal{Q}})=\sum_{i}q_{i}\text{Tr}(\varrho_{i}O_{\mathcal{Q}})=\langle O_{\mathcal{Q}}\rangle_{t_{2}}.$ (12)
It is not difficult to find that the evolution from $t_{0}$ to $t_{n}$ still has $\langle O_{\mathcal{Q}}\rangle_{t_{0}}=\langle O_{\mathcal{Q}}\rangle_{t_{n}}$ (see appendixes). This states that a hybrid dynamical model of the quantum-classical system cannot make the local $\langle O_{\mathcal{Q/C}}\rangle$ change while satisfying the conservation of global $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle$.
We will now show that the interaction of two quantum systems is exchangeable up to local observables in the case of conservation laws, which is fundamentally different from the hybrid dynamics. In quantum theory, the conserved observable $O$ satisfies
$\langle[O_{\mathcal{Q}_{1}}+O_{\mathcal{Q}_{2}},U_{\mathcal{Q}_{1}\mathcal{Q}_{2}}]\rangle=0,$ (13)
where $U_{\mathcal{Q}_{1}\mathcal{Q}_{2}}$ is the unitary evolution of composite system $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$. In general, $\langle[O_{\mathcal{Q}_{i}},U_{\mathcal{Q}_{1}\mathcal{Q}_{2}}]\rangle\neq 0$, where $i=1,2$. That is the local observable $O_{\mathcal{Q}_{i}}$ is generally not a conserved quantity. Therefore, one may have
$\langle O_{\mathcal{Q}}\rangle_{t}\neq\langle O_{\mathcal{Q}}\rangle_{t+\Delta t}.$ (14)
[…]
Based on the inconsistency between the predictions of quantum-classical and full quantum dynamics, one immediately has
Corollary 2
If the local observable $\langle O_{S}\rangle$ of a quantum system $S$ is changed after interacting with an unknown matter $E$, under the conservation law of a global quantity $\langle O_{S}\rangle+\langle O_{E}\rangle$ of the global system $SE$ and assuming that the dynamics satisfies the decomposition in figure 1, then $E$ must be non-classical.
That is, subject to a specific conservation law, if a matter of unknown nature can induce a change in the corresponding local observable of a quantum system, it may imply that this unknown matter is quantumlike, ruling out a class of hybrid quantum-classical dynamics.
IV Quantumness of gravitational field
Figure 2: Momentum change of two quantum matters in the interaction of gravitational field $E$. (a) At the moment of $T=t$, the initial positions of the two stationary masses are $x_{S}$ and $x_{M}$, and the momentum are $\langle P_{\mathcal{Q/C}}\rangle=p_{S}=p_{M}=0$. (b) At the moment of $T=t+\Delta t$, the positions of masses $S$ and $M$ have changed, and their respective momentum changes in the same amount and in opposite directions while the whole satisfies the conservation of momentum.
In this section, we will argue the quantumlike nature of gravity based on two no-go theorems we proposed. Before that, we introduce the assumption of locality. Locality is defined here as any interaction between objects at a distance that is mediated through a field, rather than being instantaneous. This asserts that the gravitational interaction between objects $A$ and $B$ is mediated through a gravitational field. That is there is no direct interaction between objects $A$ and $B$.
Now we denote the gravitational field labeled as $E$, which is primarily generated by a massive object denoted as $M$. As depicted in Figure 2 (a), at $T=t_{0}$, we suppose there are two stationary quantum masses labeled as $S$ and $M$ respectively. Based on our existing observations of the universe, these masses may mutually attract to each other as a consequence of the gravitational field $E$ mediating their interaction. While the total momentum of the system remains conserved at zero, the individual momentum of system $S$ and system $M$ increases in opposite directions, as shown in Figure 2 (b).
Let us analyze how the predictions of the classical gravitational field contradict the above description of the phenomenon. Here we delve into the discussion of the gravitational field in two cases. In the first case, the gravitational field $E$ is considered to be classical. Due to the local nature of interaction, the interaction survives only in $ES$ and $EM$. Based on the no-go theorem (1 or 2 ) , and assuming momentum conservation, the gravitational field of classical nature $E$ cannot induce any change in the momentum $P$ of quantum matter $S$. In other words,
$\langle P_{S}\rangle_{t}=\langle P_{S}\rangle_{t+\Delta t},$ (15)
where $\langle P_{S}\rangle_{t}=\text{Tr}(P\varrho)$ and $\langle P_{S}\rangle_{t}=\text{Tr}(P\varrho^{\prime})$. Moving on to the second case, let’s now consider $E$ is a quantum gravitational field. In this scenario, the Hamiltonian of composed system can be expressed as $H_{SME}=H_{S}+H_{E}+H_{M}+H_{SE}+H_{EM}$. Remarkably, the Hamiltonian allows us to meet the momentum conservation criteria, ensuring that $\langle[U_{SEM},P_{S}+P_{E}+P_{M}]\rangle=0$ while $\langle[U_{SEM},P_{S/E/M}]\rangle\neq 0$, where $U_{SEM}=\text{exp}(-iH_{SEM}T/\hbar)$. That is, one may has
$\langle P_{S}\rangle_{t}\neq\langle P_{S}\rangle_{t+\Delta t}.$ (16)
Compare to the classical gravitational field, the quantum one has ability to induce changes in the momentum of quantum system for both $S$ and $M$. Consequently, if a particle is observed to change momentum after being subjected to a gravitational field, then the gravitational field must be quantum.
Let now suppose $M$ to be a huge massive object, i.e. Earth or a metal sphere, representing quantum matter that adheres to the classical limit, as shown in Figure 2. According to our argument, under the principle of momentum conservation, any momentum change in the quantum mass $S$ resulting from the influence of $M$ mediated through the gravitational field is viewed as a sufficient evidence for the quantum nature of gravity. Actually, there are lots of experimental observation to support our argument. For example, both the free fall of a quantum particle and the displacement of particle caused by gravitational interaction of the metal sphere [51] offer compelling indications of the quantumness of gravity.
V Other conservation laws
While we have argued for the non-classical nature of gravity in terms of conservation of momentum, a natural question is the why conservation of momentum is special? Is it possible to derive the same conclusion using other conservation laws, such as conservation of energy? The answer is yes.
When we consider the gravitational field as an interacting object, the energy of the particle may be just its own free energy and may not include the gravitational potential energy. Imagine a particle’s energy, defined as free energy, i. e. kinetic energy. Our above analysis asserts that if the kinetic energy of a particle increases following an interaction with a gravitational field, this significant observation implies that gravity behaves in a manner inconsistent with classical objects, assuming the principle of energy conservation holds true [52]. Evidently, the free-fall of a particle may provide compelling evidence for the quantum nature of gravity .
VI Conclusions
We have discussed two general framework for quantum-classical hybrid dynamics, including Hamiltonian formalism and an operational one, offering an approach that is model-independent. Both frameworks are predicated on the fundamental premise that any classical system can only has one basis and maintain classical correlations with a quantum system.
Within the frameworks, we present no-go theorems that pertains to systems subject to a conservation law. The no-go theorems explore a fundamental limitation of classical system to interact with a quantum system, assuming that the interaction satisfies Hamiltonian formalism or the decomposition property set out in figure 1, i. e. any classical system can not induce a change of the local observable of quantum system while a global conserved quantity of hybrid system is assumed. Based on these no-go theroems, our exploration delves into the quantum nature of gravity, unveiling a profound connection that exists between conservation laws and the inherent quantum aspects of gravity.
Additionally, our analysis underscores the pivotal role played by expectation values in facilitating a holistic comprehension of the physical universe. This transcends the limitations imposed by disparate mathematical structures and serves as a bridge, connecting various physical theories into a unified description.
While our argument regarding the quantum nature of gravity remains within a non-relativistic context, its generality and validity persist. We expect to discover more relationships between fundamental principles and gravity, leading to a deeper understanding of the nature of spacetime.
Acknowledgements.
T. F. thanks the support of QICI of the University of Hong Kong, and is grateful to Tian Zhang for helpful discussions. C. M. and V.V. thank the Gordon and Betty Moore Foundation. This publication was also made possible through the support of the ID 61466 grant from the John Templeton Foundation, as part of the The Quantum Information Structure of Spacetime (QISS) Project (qiss.fr). The opinions expressed in this publication are those of the Authors and do not necessarily reflect the views of the John Templeton Foundation.
Appendix A Observable in general quantum-classical hybrid systems
Different physical theories, such as classical mechanics and quantum mechanics, may exhibit incompatible mathematical structures. However, they share a common feature: the ability to describe physical systems using expectation values. In terms of expectation values, both classical and quantum theories yield real numbers. The validity of a theory is determined by whether its predicted expectations match those observed experimentally. Therefore, expectation values of observables (dynamic variables) serve as a bridge connecting different theories. From a measurable perspective, a general description of expectation value of observable is required.
In the real physical world, all we have access to are the expectation values of observables (dynamics variables). And these expectation values are jointly determined by a physical state and observables. That is, state and observables determine the attributes (expectation values of observables) of a physical system. Here we describe them as
[…]
In quantum theory, the expectation value of an observable $O$ is given as
$\langle O\rangle=\{\varrho,O\}=\text{Tr}(\varrho{O}),$ (18)
where $\varrho$ is the quantum state. In classical theory, if the physical system is in a mixture of several states of different pure $\text{State}_{i}$ with probability $p_{i}$, this mixed classical state is denoted as $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$. Using the language defined above, the expectation value of a classical system can be described as
$\{\wp_{\mathcal{C}},O\}\equiv\sum_{i}p_{i}\{\text{State}_{i},O\}.$ (19)
As can be seen, even if we do not know what the mechanism of the classical sector is, one can still use this abstraction to represent the expectation value of an observable $O$. Note that the classical state $\wp_{\mathcal{C}}$ contain information about an infinite number of dynamical variables. Above we are only concerned with one observable.
Now let’s define the evolution of observables in general. From the initial moment ($t_{0}$) to the next moment ($t_{1}$), the evolution of the expectation value of is described as
[…]
Obviously, it indicates $\langle O\rangle_{t_{0}}\rightarrow\langle O\rangle_{t_{1}}$. In the above description, we define the state as varying with time. In fact one can also define the case where the observable varies with time and the state is fixed.
Let us now analyse how hybrid systems are described in this framework. A hybrid classical-quantum system is describe as $\mathcal{C}\oplus\mathcal{Q}$, where $\mathcal{C}$ and $\mathcal{Q}$ represent classical and quantum sectors respectively. The expectation value of an (additive) observable is given
{Hybrid $\{\text{Hybrid State},O\}=\{\text{Hybrid State},O_{\mathcal{C}}+O_{\mathcal{Q}}\},$ (21)
where $O_{\mathcal{C}}$ and $O_{\mathcal{Q}}$ are the observable of classical and quantum systems respectively. Since it is impossible for the quantum and classical sectors of a hybrid system to be quantum mechanically correlated (without any quantum correlation), the most general correlation that can exist for a hybrid system is the classical correlation. Therefore, one has
{Hybrid $\{\text{Hybrid State},O\}=\{\text{Hybrid State},O_{\mathcal{C}}+O_{\mathcal{Q}}\}=\langle O_{\mathcal{C}}\rangle+\langle O_{\mathcal{Q}}\rangle,$ (22)
where
$\begin{split}\langle O_{\mathcal{C}}\rangle&=\{\text{ Reduced Classical State},O_{\mathcal{C}}\},\\ \langle O_{\mathcal{Q}}\rangle&=\{\text{ Reduced Quantum State},O_{\mathcal{Q}}\}.\end{split}$ (23)
That is $O_{\mathcal{C}}$ and $O_{\mathcal{Q}}$ are the expected values of the local systems. Generally, all mixed states can be categorized into two cases, one in which there is no classical correlation between the classical and quantum sectors, and the other in which there is a classical correlation. Suppose the initial state of hybrid system $\chi_{\mathcal{QC}}$ without classical correlation, that is $\chi_{\mathcal{QC}}=\wp_{\mathcal{C}}\oplus\varrho$, then
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle=\{\wp_{\mathcal{C}}\oplus\varrho,O_{\mathcal{C}}+O_{\mathcal{Q}}\}=\{\wp_{\mathcal{C}},O_{\mathcal{C}}\}+\{\varrho,O_{\mathcal{Q}}\}=\langle O_{\mathcal{C}}\rangle+\langle O_{\mathcal{Q}}\rangle.$ (24)
[…]
$\{\sum_{j}q_{j}\wp_{\mathcal{C}j}\oplus\varrho_{j},O_{\mathcal{C}}+O_{\mathcal{Q}}\}=\sum_{j}q_{j}\{\sum_{i}p^{(j)}_{i}\text{State}^{(j)}_{i},O_{\mathcal{C}}\}+\sum_{j}q_{j}\text{Tr}(\varrho_{j}O_{\mathcal{Q}})=\langle O_{\mathcal{C}}\rangle+\langle O_{\mathcal{Q}}\rangle,$ (25)
where $\wp_{\mathcal{C}j}=\sum_{i}p^{(j)}_{i}\text{State}^{(j)}_{i}$. In general, the the evolutionary final state of the hybrid state $\wp_{\mathcal{C}}\oplus\varrho$ (no classical correlation) generally contains classical correlations, and the corresponding evolution of the expectation value is
$\{\wp_{\mathcal{C}}\oplus\varrho,O_{\mathcal{C}}+O_{\mathcal{Q}}\}\rightarrow\{V\cdot\wp_{\mathcal{C}}\oplus\varrho,O_{\mathcal{C}}+O_{\mathcal{Q}}\}\rightarrow\{\sum_{i}q_{i}\text{State}_{i}\oplus\Lambda_{i}\cdot\varrho,O_{\mathcal{C}}+O_{\mathcal{Q}}\}=\langle O_{\mathcal{C}}\rangle+\langle O_{\mathcal{Q}}\rangle$ (26)
where $V$ and $\Lambda_{i}$ represent classical and quantum maps respectively and satisfying $V\cdot\wp_{\mathcal{C}}=\sum_{i}q_{i}\text{State}_{i}$ and $\Lambda_{i}\cdot\varrho=\varrho_{i}$. One may see that the changes of local observables $\langle O_{\mathcal{C}}\rangle$ and $\langle O_{\mathcal{Q}}\rangle$ may indicate the backaction of the quantum system to the classical system, and vice versa.
Appendix B Multiple-time evolution of hybrid systems
In the main text, we have analysed the dynamics of the hybrid system for an infinitesimal time interval. It is shown that for a globally additive conserved quantity, any quantum-classical hybrid dynamics is fundamentally incapable of varying the observables (corresponding to the conserved quantity) in either subsystem.
Here, we analyse in more detail the dynamics and properties of the hybrid system in terms of its evolution over multiple time intervals. As shown in Fig. 3, the general circuit of the evolution (from $t_{0}$ to $t_{2n}$) of the hybrid system is presented. Similar to the analyzed in main text, the composed state of the hybrid system at $t_{2m}$ is donated as
$\chi_{\mathcal{QC}}(t_{2m})=\sum_{i}q^{(2m)}_{i}\wp^{(2m)}_{\mathcal{C}i}\oplus\varrho^{(2m)}_{i},$ (27)
where $q^{(2m)}_{i}$ leads to classical correlation between quantum and classical sectors. It is worth noting that odd time moments correspond to $V$ operations which is locally act on classical system, while even ones correspond to $\Lambda$ operations acting on whole systems, as shown in Fig. 3. Therefore, $\varrho_{i}^{(2m-2)}=\varrho_{i}^{(2m-1)}$.
Suppose $\wp^{(2m-1)}_{\mathcal{C}i}=\sum_{j}p^{(2m-1)}_{ij}\text{State}_{j}$, then the state of hybrid system evolve from $t_{2m-1}$ (odd) to $t_{2m}$ (even) is given as
$\chi_{\mathcal{QC}}(t_{2m-1})=\sum_{i}q^{(2m-1)}_{i}\wp^{(2m-1)}_{\mathcal{C}i}\oplus\varrho^{(2m-1)}_{i}\rightarrow\chi_{\mathcal{QC}}(t_{2m})=\sum_{i,j}p^{(2m-1)}_{ij}q^{(2m-1)}_{i}\text{State}_{j}\oplus\Lambda_{j}^{(2m)}\cdot\varrho^{(2m-1)}_{i}.$ (28)
That is
$\chi_{\mathcal{QC}}(t_{2m})=\sum_{i}q^{(2m)}_{i}\wp^{(2m)}_{\mathcal{C}i}\oplus\varrho^{(2m)}_{i}=\sum_{i,j}p^{(2m-1)}_{ij}q^{(2m-1)}_{i}\text{State}_{j}\oplus\Lambda_{j}^{(2m)}\cdot\varrho^{(2m-1)}_{i}.$ (29)
Similarly, one can analyze the nature of the dynamics from $t_{2m-2}$(even) to $t_{2m-1}$ (odd) instances,
$\chi_{\mathcal{QC}}(t_{2m-2})=\sum_{i}q^{(2m-2)}_{i}\wp^{(2m-2)}_{\mathcal{C}i}\oplus\varrho^{(2m-2)}_{i}\rightarrow\chi_{\mathcal{QC}}(t_{2m-1})=\sum_{i}q^{(2m-2)}_{i}V^{2m-1}\cdot\wp^{(2m-2)}_{\mathcal{C}i}\oplus\varrho^{(2m-2)}_{i}.$ (30)
Figure 3: Quantum-classical dynamics of multi-time evolution. Each time interval is infinitesimal, that is, $t_{i}-t_{i+1}=\delta t$. The dynamics of a hybrid system consists of these infinitesimal slices of evolution.
Now let assume that there is a conserved observable $O_{\mathcal{C}}+O_{\mathcal{Q}}$. The expectation value of $O_{\mathcal{C}}+O_{\mathcal{Q}}$ at time $t_{2m-1}$ and $t_{2m}$ are given as
$\begin{split}\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m-1}}&=\{\chi_{\mathcal{QC}}(t_{2m-1}),O_{\mathcal{C}}\}+\{\chi_{\mathcal{QC}}(t_{2m-1}),O_{\mathcal{Q}}\}\\ &=\sum_{j}q^{(2m-1)}_{i}\{\wp^{(2m-1)}_{\mathcal{C}i},O_{\mathcal{C}}\}+\sum_{i}q^{(2m-1)}_{i}\text{Tr}(\varrho^{(2m-1)}_{i}O_{\mathcal{Q}})\end{split}$ (31)
and
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m}}=\{\chi_{\mathcal{QC}}(t_{2m}),O_{\mathcal{C}}\}+\{\chi_{\mathcal{QC}}(t_{2m}),O_{\mathcal{Q}}\}=\sum_{j}q^{(2m)}_{i}\{\wp^{(2m)}_{\mathcal{C}i},O_{\mathcal{C}}\}+\sum_{i}q^{(2m)}_{i}\text{Tr}(\varrho^{(2m)}_{i}O_{\mathcal{Q}}),$ (32)
respectively. Substituting Eq. (29) into Eq. (32), it is easy to obtain $\{\chi_{\mathcal{QC}}(t_{2m-1}),O_{\mathcal{C}}\}=\{\chi_{\mathcal{QC}}(t_{2m}),O_{\mathcal{C}}\}$. Thus $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m}}$ can rewritten as
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m}}=\{\chi_{\mathcal{QC}}(t_{2m}),O_{\mathcal{C}}\}+\{\chi_{\mathcal{QC}}(t_{2m}),O_{\mathcal{Q}}\}=\sum_{j}q^{(2m-1)}_{i}\{\wp^{(2m-1)}_{\mathcal{C}i},O_{\mathcal{C}}\}+\sum_{i}q^{(2m)}_{i}\text{Tr}(\varrho^{(2m)}_{i}O_{\mathcal{Q}}),$ (33)
Since $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m-1}}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m}}$, one obtains
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m-1}}-\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m}}=\sum_{i}q^{(2m-1)}_{i}\text{Tr}(\varrho^{(2m-1)}_{i}O_{\mathcal{Q}})-\sum_{i}q^{(2m)}_{i}\text{Tr}(\varrho^{(2m)}_{i}O_{\mathcal{Q}})=0.$ (34)
In other word, observable $O_{\mathcal{Q}}$ is conserved from odd to even time moments
$\langle O_{\mathcal{Q}}\rangle_{t_{2m-1}}=\langle O_{\mathcal{Q}}\rangle_{t_{2m}}.$ (35)
Similarly, the state of hybrid system evolve from $t_{2m-2}$ to $t_{2m-1}$ is given as
$\chi_{\mathcal{QC}}(t_{2m-2})=\sum_{i}q^{(2m-2)}_{i}\wp^{(2m-2)}_{\mathcal{C}i}\oplus\varrho^{(2m-2)}_{i}\rightarrow\chi_{\mathcal{QC}}(t_{2m-1})=\sum_{i}q^{(2m-2)}_{i}V^{2m-1}\cdot\wp^{(2m-2)}_{\mathcal{C}i}\oplus\varrho^{(2m-2)}_{i}.$ (36)
Therefore, the state at $t_{2m-1}$ is rewritten as
$\chi_{\mathcal{QC}}(t_{2m-1})=\sum_{i}q^{(2m-1)}_{i}\wp^{(2m-2)}_{\mathcal{C}i}\oplus\varrho^{(2m-1)}_{i}=\sum_{i}q^{(2m-2)}_{i}V^{(2m-1)}\cdot\wp^{(2m-1)}_{\mathcal{C}i}\oplus\varrho^{(2m-2)}_{i}.$ (37)
Immediately, the observable $O_{\mathcal{C}}+O_{\mathcal{Q}}$ for $t_{2m-2}$ and $t_{2m-1}$ are
$\begin{split}\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m-2}}=\{\chi_{\mathcal{QC}}(t_{2m-2}),O_{\mathcal{C}}+O_{\mathcal{Q}}\}=\sum_{j}q^{(2m-2)}_{i}\{\wp^{(2m-2)}_{\mathcal{C}i},O_{\mathcal{C}}\}+\sum_{i}q^{(2m-2)}_{i}\text{Tr}(\varrho^{(2m-2)}_{i}O_{\mathcal{Q}}),\\ \langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2m-1}}=\{\chi_{\mathcal{QC}}(t_{2m-1}),O_{\mathcal{C}}+O_{\mathcal{Q}}\}=\sum_{j}q^{(2m-2)}_{i}\{V^{(2m-1)}\cdot\wp^{(2m-2)}_{\mathcal{C}i},O_{\mathcal{C}}\}+\sum_{i}q^{(2m-2)}_{i}\text{Tr}(\varrho^{(2m-2)}_{i}O_{\mathcal{Q}}),\\ \end{split}$ (38)
where $\{\chi_{\mathcal{QC}}(t_{j}),O_{\mathcal{C}}+O_{\mathcal{Q}}\}=\{\chi_{\mathcal{QC}}(t_{j}),O_{\mathcal{C}}\}+\{\chi_{\mathcal{QC}}(t_{j}),O_{\mathcal{Q}}\}$.
Obviously, from $t_{2m-2}$ to $t_{2m-1}$, there is no action on quantum sector, $\{\chi_{\mathcal{QC}}(t_{2m-2}),O_{\mathcal{Q}}\}=\{\chi_{\mathcal{QC}}(t_{2m-1}),O_{\mathcal{Q}}\}$. Namely,
$\langle O_{\mathcal{Q}}\rangle_{t_{2m-2}}=\langle O_{\mathcal{Q}}\rangle_{t_{2m-1}}.$ (39)
Thus, for a conserved observable $O_{\mathcal{C}}+O_{\mathcal{Q}}$, at any $t_{m}$ and $t_{n}$,
$\langle O_{\mathcal{Q}}\rangle_{t_{m}}=\langle O_{\mathcal{Q}}\rangle_{t_{n}},\langle O_{\mathcal{C}}\rangle_{t_{m}}=\langle O_{\mathcal{C}}\rangle_{t_{n}}.$ (40)
It means that the expectation value of observable $O_{\mathcal{Q/C}}$ of the subsystem is also a conserved quantity in the quantum-classical hybrid dynamics. Therefore, our no-go theorem applies to the evolution of hybrid systems at any time.