### v4

# Conservation Laws and the Quantization of Gravity

Abstract
Adopting general frameworks for quantum-classical dynamics, we analyze the interaction between quantum matter and a classical gravitational field. We point out that, assuming the conservation of momentum or energy, and assuming that the dynamics obey a general decomposition property, the classical gravitational field cannot change the momentum or energy of the quantum system, whereas the quantum gravitational field can do so. Drawing upon the fundamental relationship between conservation laws and the quantum properties of objects, our analysis offers new perspectives for the study of quantum gravity and provides a novel interpretation of existing experimental observations, such as free fall.
The quantization of gravity, a fundamental issue of modern theoretical physics, has captivated the imaginations of physicists for over a century. This profound quest seeks to reconcile two pillars of contemporary physics: general relativity and quantum theory. Yet, it is hard to unify them into a unique corpus of laws. Several approaches, such as field theory, “geometrization” of quantum mechanics and string theory, have emerged in the pursuit of reconciling quantum theory with general relativity [1, 2, 3]. However, all these approaches face significant theoretical and conceptual challenges [4, 5], which have prompted some to propose the fundamental classical nature of gravity [6]. Further, some have even argued that quantizing gravity is not needed after all [7] or that gravity may not even be a fundamental force [8, 9]. Whether or not the quantisation of gravity is necessary has become highly contentious [10, 11, 12, 13, 15, 14], primarily due to the enduring difficulty in acquiring direct supporting evidence [16, 17].
As observed by Feynman during the 1950 Chapel Hill conference, it was pointed out that the physical state of the composite system exhibits distinct properties depending on whether the mass is entangled with gravity or merely classically correlated with it [20]. This highlights the importance of the superposition of large masses in the study of quantum gravity. Recently, there has been a revival of Feynman’s idea , which highlights gravitationally induced entanglement between two massive particles providing strong evidence on whether the gravitational field is mediated by the transfer of quantum information [18, 19]. This is clearly sufficient evidence for quantum effects in gravity. Unfortunately, gravitationally induced entanglement is extremely weak, and even though there are some advanced theoretical analyses [21, 22, 23, 24] and experimental protocols [25, 26, 27, 28, 29, 30] available, with the current experimental techniques and advances, it is not yet time to observe this effect.
The progress outlined above based on the quantum information perspective is truly enlightening, even though the pursuit of quantum entanglement induced by gravity remains a formidable challenge. A natural question arises: Are the existing experimental observations of gravity sufficient to analyze the quantum nature of gravity? In this paper, we unequivocally affirm this possibility.
In historical context, Eppley and Hannah endeavored to construct a logical argument challenging the classical model of gravity’s compatibility with the conservation of momentum [10]. It is essential to acknowledge that their work predominantly consisted of a thought experiment and lacked empirical observations. Subsequently, their assertions have faced widespread skepticism and have been refuted by subsequent research [11, 13]. Moreover, other gedankenexperiments have been conducted, highlighting the perplexing paradoxes of causality and complementarity arising from classical gravity [31, 32]. Indeed, the introduction of some basic principles and physical laws helps us to analyse the non-classical nature of matter, even in the absence of entanglement witnesses [33, 34, 35, 36].
Inspired by [34, 35], our work aims to delve into the intrinsic relationship between conservation laws and the underlying nature of gravity. We shall aim at relaxing some of the assumptions in [34], e.g. linearity of the overall dynamics, by adding another constraint on the dynamics, outlined in section II. To accomplish this, we harness information-theoretic methodologies to construct a general framework of hybrid systems that amalgamates quantum and classical dynamics into a unified whole [37]. Within this framework, we present a no-go theorem, establishing a fundamental limitation. It stipulates that any quantum-classical dynamics cannot induce alterations in the local observables of either classical or quantum systems, all the while maintaining the conservation of global observables. Our results bring to light a profound revelation: when considering the conservation of either momentum or energy, the classical gravitational field is unable to effect changes in the momentum or energy of a quantum system. In other words, in this frameworks, conservation laws are sufficient but not necessary conditions for the quantization of gravity. This proposition stands in contrast to established experimental observations, such as the free fall experiment. Conversely, our arguments compellingly demonstrate that the quantum gravitational field possesses the capability to enact such changes, offering compelling evidence for the quantum nature of gravity.
I Definition of a Classical system
A fundamental distinction between classical and quantum systems lies in their contextual nature. In this paper, by “classical” we mean precisely a system that has only one single observable (single basis, e.g. $\{|i\rangle\langle i|\}$). The other observables of a classical system are functions of this observable [36, 14]. The definition of a classical system here does not only require that its corresponding physical states are all perfectly distinguishable, but it also implies that it is restricted from interacting with any matter, i.e., only these commutative observables are involved in the interaction [38, 19, 39].
A natural consequence of the definition of classical systems is that classical systems are only classically correlated with any system, including quantum systems [40]. Since a system without two or more than two non-commutative observables is not able to establish a quantum correlation with a quantum system. Besides, one may infer that a classical system can perform different control operations on quantum systems, while quantum systems cannot [41].
II General Quantum-classical Dynamics
Classical correlations manifest in the majority of physical dynamic processes. These correlations can arise between quantum-quantum systems, classical-classical systems, or quantum-classical systems. In the subsequent discussion, our attention will be directed towards exploring general quantum-classical hybrid systems. We emphasize that the Hamiltonian formalism of closed systems, i.e., the unitary quantum theory, is a special case in this framework (see Appendix). Before proceeding, we will provide the necessary notations and definitions for both quantum and classical systems.
For a quantum system, we utilize the density matrix in Hilbert space to represent a quantum state. On the other hand, for a classical system, we do not make any assumptions about a specific model. To maintain generality, we denote a pure classical state as $\text{State}_{i}$”, where each unique value of $i$ represents a perfectly distinguishable state. Naturally, for a mixed classical state, it can be defined as follows: $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$, where $p_{i}$ is the probability of different pure states $\text{State}_{i}$” (for continuous classical system, $\wp_{\mathcal{C}}=\int_{x}p(x)\text{State}_{x}dx$). In the following, without loss of generality, we focus on the discrete case, as the continuous case can be approximated by the discrete case with arbitrarily large dimensions and arbitrary precision.
As classical systems cannot be quantum correlated with quantum systems, their composite states exhibit solely classical correlations. Suppose the initial state of the hybrid system lacks classical correlation, following a quantum-classical interaction, the resulting state generally evolves and can be expressed in the following general form:
$\wp_{\mathcal{C}}\oplus\varrho\rightarrow\sum_{i}q_{i}\wp_{{\mathcal{C}}i}\oplus\varrho_{i},$ (1)
where $\{q_{i}\}$ is the probability distribution of classical correlation and $\wp_{{\mathcal{C}}i}=\sum_{j}p_{j}\text{State}_{i}$. Historically, the evolutionary state of a hybrid system is called the hybrid density operator [42, 43], which satisfies an overall normalization condition and preserves the form of the state.
An operational definition of quantum-classical dynamics $\Theta_{\mathcal{QC}}$ can be derived based on the limitations of classical correlations and the properties of classical systems. As depicted in Figure 1, we provide a general operational definition for quantum-classical dynamics over infinitesimal time intervals. In this context, we refrain from imposing any specific theoretical framework on quantum-classical dynamics and instead focus on its operational decomposition, emphasizing how it generates classical correlations. This decomposition aligns with the principles of physical causality. In non-relativistic quantum mechanics, the Hamiltonian dictates that the evolution of interacting particles can occur simultaneously, i.e. $e^{-i(\sum_{j}H_{j}+\sum_{k}H^{k}_{int})t}$. However, in reality, interactions between atoms are mediated by electromagnetic fields, propagating at the finite speed of light. Consequently, the actual interaction processes are sequential rather than simultaneous, akin to Trotterization, i.e. $(\prod_{j}e^{-iH_{j}\delta t}\prod_{k}e^{-iH^{k}_{int}\delta t})^{t/{\delta t}}$. Furthermore, in the linear version of the Diósi-Penrose model, the ultimate master equation is also given by a sequential evolution [44]. This consists of a free evolution governed by the stochastic master equation for monitoring the density of masses, followed by the gravitational potential acting over an infinitesimal amount of time [44]. Thus, even if whole operations may appear consistent, the sequential decomposition and whole operations embody distinct physical meanings. The global viewpoint may lose the underlying physics.
Figure 1: Operational definition of general quantum-classical dynamics. Quantum-classical dynamics of multi-time evolution: Each time interval is infinitesimal, that is, $t_{i}-t_{i+1}=dt$. The dynamics of a hybrid system consist of sequentially infinitesimal slices of local evolution and global evolution. This decomposition satisfies the causality.
Without loss of generality, here we analyze the hybrid dynamics of an infinitesimal time slice, as in figure 1, while the proof of continuous time is put into the Appendix. We suppose the initial state of the hybrid system is uncorrelated, denoted as $\chi_{\mathcal{QC}}=\wp_{\mathcal{C}}\oplus\varrho$, where $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$. Generally, as shown in Figure 1 (single segment), one may divide the quantum-classical dynamics into two distinct parts. Firstly, there is a map that is physically and effectively equivalent to the local action $V=V_{\mathcal{C}}\oplus V_{\mathcal{Q}}$ where $V_{\mathcal{Q/C}}$ operating on the quantum /classical state $\varrho_{\mathcal{Q}}/\wp_{\mathcal{C}}$. Consequently, $V$ transforms the hybrid state $\chi_{\mathcal{QC}}$ into $V(\wp_{\mathcal{C}}\oplus\varrho)=V_{\mathcal{C}}(\wp_{\mathcal{C}})\oplus V_{\mathcal{Q}}(\varrho)=\sum_{i}q_{i}\text{State}_{i}\oplus\varrho^{\prime}$. The second part involves the establishment of a control channel $\Lambda_{i}$ from the classical sector to the quantum sector, from which the classical correlations of the states of the composite system originate. The specific process of the second part can be expressed as:
$\sum_{i}q_{i}\text{State}_{i}\oplus\varrho^{\prime}\rightarrow\sum_{i}q_{i}\text{State}_{i}\oplus\varrho_{i},$ (2)
where $\rho_{i}=\Lambda_{i}(\varrho)$. Therefore, one has $\Theta_{\mathcal{QC}}(\chi_{\mathcal{QC}})=\chi_{\mathcal{QC}}^{\prime}=\sum_{i}q_{i}\text{State}_{i}\oplus\varrho_{i}$. Note that the dynamics $\Theta_{\mathcal{QC}}$, $V$ and $\Lambda$ don’t need to be (input) state-independent local operations and control operations, but rather state-dependent. Since for a general hybrid dynamics $\Theta_{\mathcal{QC}}$ one does not exclude the possibility that it is a nonlinear mapping (no linearity and reversibility need to be assumed). It doesn’t matter what the nature of the hybrid dynamics is, as long as it ends up embodying the state of classical correlation. Thus, in general, $\Theta_{\mathcal{QC}},V$, $\Lambda_{i}$ and $q_{i})$ are may be the function of $\chi_{\mathcal{QC}}$. For convenience, we abbreviate them as $V$, $\Lambda_{i}$, and $q_{i}$ below.
The above operational definition of a hybrid system requires that a legal physical state is mapped to another physical state at any moment [42, 43]. An alternative approach to comprehend this decomposition is to liken the classical state $\{\text{State}_{i}\}$ to a set of orthogonal bases $\{|i\rangle\langle i|\}$. This analogy arises from the fact that the classical system only possesses one basis, and any arbitrary mapping of its interactions with the quantum system can always be broken down into local operations and control operations of the classical to the quantum system [50]. Our framework can be reduced in the quantum theory case if the linearity holds. For example, in quantum theory, $e^{-i|j\rangle\langle j|\otimes A_{j}}$ can be exactly represented by the control operation $\Lambda$ with unitary $\Lambda_{i}$. As we mentioned before, instead of assuming linearity, we assume that the control of a classical system by a quantum system is not allowed, hence, the most general form of quantum-classical dynamics over infinitesimal time intervals must be one that cascades from the circuit shown in Figure 1.
II.1 Observables in Hybrid systems
The attributes of a physical system, specifically the expectation values of observables, are jointly determined by the state and observables. Here the relationship between the state of a physical system and its observable can be succinctly described as
Observable{State, $\langle\text{Observable}\rangle\equiv\{\text${State, Observable}\}. (3)
Both classical and quantum theories can be encompassed within this language. In quantum theory, the expectation value of an observable $O$ is given as $\{\varrho,O\}\equiv\text{Tr}(\varrho{O})$, where $\varrho$ is the quantum state (note that here $\{,\}$ is not an anti-commutative operator). Without loss of generality, for simplicity, we only analyze the case where the classical system is discrete.
In a classical theory, the expectation value of a classical system can be described as $\{\wp_{\mathcal{C}},O\}\equiv\sum_{i}p_{i}\{\text{State}_{i},O\}$, where $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$. In particular, if the $\text{State}_{i}$ is the eigenstate of eigenvalue $O_{i}$, then $\sum_{i}p_{i}\{\text{State}_{i},O\}=\sum_{i}p_{i}O_{i}$, which is similar to the definition of legitimate expectation value $\langle O\rangle$. For a hybrid quantum-classical system, the expectation value of an observable is naturally defined as follows $\langle O\rangle=\{\chi_{\mathcal{QC}},O\},$ where $\chi_{\mathcal{QC}}$ is the quantum-classical hybrid state defined above. If $O$ is an additive observable, i.e. $O=O_{\mathcal{C}}+O_{\mathcal{Q}}$, $O_{\mathcal{C}}+O_{\mathcal{Q}}=\{\chi_{\mathcal{QC}},O_{\mathcal{C}}\}+\{\chi_{\mathcal{QC}},O_{\mathcal{Q}}\}$ ( See Appendixes for more details of observables in hybrid systems).
II.2 A No-go theorem
Here we present a no-go theorem that demonstrates the impossibility of a general quantum-classical dynamics (satisfying the decomposition in figure 1) causing a change in the local observable of either a classical or quantum system while preserving the conservation of the global observable. In other words, the conservation law imposes limitations on the exchange of corresponding local observables between classical and quantum systems. The no-go theorem is as follows:
Theorem 1
Any quantum-classical dynamics that satisfies the decomposition in figure 1 can not change the local observable $\langle O_{\mathcal{C/Q}}\rangle$ of either a classical or quantum system under the conservation law of a global quantity of the hybrid system $\mathcal{C}\oplus\mathcal{Q}$.
Let’s see how this theorem is derived. Suppose there is a universal conservation law for the observable $O$, such as energy or momentum conservation. For any physical system ($\mathcal{C}$ or $\mathcal{Q}$) without interaction with others, the observable $O_{\mathcal{C}/\mathcal{Q}}$ should remain unchanged, as the universal conservation law holds. That is, the free evolution of a system does not change $O_{\mathcal{C}/\mathcal{Q}}$. For a composed system of $\mathcal{C}$ and $\mathcal{Q}$ (a quantum-classical hybrid system), the conserved observable should be $O=O_{\mathcal{C}}+O_{\mathcal{Q}}$, satisfying
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t+\Delta t},$ (4)
where $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle=\{\chi_{\mathcal{QC}},O_{\mathcal{C}}\}+\{\chi_{\mathcal{QC}},O_{\mathcal{Q}}\}=\langle O_{\mathcal{C}}\rangle+\langle O_{\mathcal{Q}}\rangle$. The conservation law dictates that equation (4) must hold at all times. This implies that, as illustrated in Figure 1, every moment of the dynamical evolution of the hybrid system must adhere to the conservation law. In other words, any local operation and the global operation must independently comply with the conservation of local/global observable $O$, respectively.
Without loss generality, suppose the initial state of the hybrid system has no classical correlation, that is $\chi_{\mathcal{QC}}=\wp_{\mathcal{C}}\oplus\varrho$, where $\wp_{\mathcal{C}}=\sum_{i}p_{i}\text{State}_{i}$. Immediately, one may calculate the corresponding observable $O$ as
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{0}}=\sum_{i}p_{i}\{\text{State}_{i},O_{\mathcal{C}}\}+\text{Tr}(\varrho O_{\mathcal{Q}}).$ (5)
As we have analyzed before, after the operation $V$, the hybrid state become $V(\wp_{\mathcal{C}}\oplus\varrho)=V_{\mathcal{C}}(\wp_{\mathcal{C}})\oplus V_{\mathcal{Q}}(\varrho)=\sum_{i}q_{i}\text{State}_{i}\oplus\varrho^{\prime}$, one has
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{1}}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}+\text{Tr}(\varrho^{\prime}O_{\mathcal{Q}}).$ (6)
Since $V_{\mathcal{C/Q}}$ is local free evolution and according to the conservation law, we have $\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}$ and $\text{Tr}(\varrho O_{\mathcal{Q}}=\text{Tr}(\varrho^{\prime}O_{\mathcal{Q}})$. Similarly, the control operation $\Lambda$ in Figure 1 changes the state from $\sum_{i}q_{i}\text{State}_{i}\oplus\varrho$ to $\sum_{i}q_{i}\text{State}_{i}\oplus\varrho_{i}$ with
$\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2}}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}+\sum_{i}q_{i}\text{Tr}(\varrho_{i}O_{\mathcal{Q}}),$ (7)
where $\rho_{i}=\Lambda_{i}(\varrho^{\prime})$. Since $O$ is conserved at any time, there must be $\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{0}}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{1}}=\langle O_{\mathcal{C}}+O_{\mathcal{Q}}\rangle_{t_{2}}$. Since $\sum_{i}p_{i}\{\text{State}_{i},O_{\mathcal{C}}\}=\sum_{i}q_{i}\{\text{State}_{i},O_{\mathcal{C}}\}$, we have
$\langle O_{\mathcal{Q}}\rangle_{t_{0}}=\text{Tr}(\varrho O_{\mathcal{Q}})=\text{Tr}(\varrho^{\prime}O_{\mathcal{Q}})=\sum_{i}q_{i}\text{Tr}(\varrho_{i}O_{\mathcal{Q}}).$ (8)
That is $\langle O_{\mathcal{Q}}\rangle_{t_{0}}=\langle O_{\mathcal{Q}}\rangle_{t_{1}}=\langle O_{\mathcal{Q}}\rangle_{t_{2}}$. It is not difficult to find that the evolution from $t_{0}$ to $t_{n}$ still has $\langle O_{\mathcal{Q}}\rangle_{t_{0}}=\langle O_{\mathcal{Q}}\rangle_{t_{n}}$ (see appendixes). This states that a hybrid dynamical model of the quantum-classical system cannot make the local $\langle O_{\mathcal{Q/C}}\rangle$ change while satisfying the conservation law of $O$.
We will now show that the interaction of two quantum systems is exchangeable up to local observables in the case of conservation laws, which is fundamentally different from the hybrid dynamics. In quantum theory, the conserved observable $O$ satisfies
$\langle[O_{\mathcal{Q}_{1}}+O_{\mathcal{Q}_{2}},U_{\mathcal{Q}_{1}\mathcal{Q}_{2}}]\rangle=0,$ (9)
where $U_{\mathcal{Q}_{1}\mathcal{Q}_{2}}$ is the unitary evolution of composite system $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$. In general, $\langle[O_{\mathcal{Q}_{i}},U_{\mathcal{Q}_{1}\mathcal{Q}_{2}}]\rangle\neq 0$, where $i=1,2$. That is, the local observable $O_{\mathcal{Q}_{i}}$ is generally not a conserved quantity. Therefore, one may have
$\langle O_{\mathcal{Q}}\rangle_{t}\neq\langle O_{\mathcal{Q}}\rangle_{t+\Delta t}.$ (10)
Note that the above theorem also holds for a continuous classical system, i.e., the classical system $\wp_{\mathcal{C}}=\int_{x}p(x)\text{State}_{x}dx$.
Based on the inconsistency between the predictions of quantum-classical and full quantum dynamics, one immediately has
Corollary 1
If the local observable $\langle O_{S}\rangle$ of a quantum system $S$ is changed after interacting with an unknown matter $E$, under the conservation law of a global quantity $\langle O_{S}\rangle+\langle O_{E}\rangle$ of the global system $SE$ and assuming that the dynamics satisfies the decomposition in figure 1, then $E$ must be non-classical.
That is, subject to a specific conservation law if a matter of unknown nature can induce a change in the corresponding local observable of a quantum system, it may imply that this unknown matter is quantumlike, ruling out a class of hybrid quantum-classical dynamics.
III Quantumness of gravitational field
Figure 2: Momentum change of two quantum matters in the interaction of gravitational field $E$. (a) At the moment of $T=t$, the initial positions of the two stationary masses are $x_{S}$ and $x_{M}$, and the momentum are $\langle P_{\mathcal{Q/C}}\rangle=p_{S}=p_{M}=0$. (b) At the moment of $T=t+\Delta t$, the positions of masses $S$ and $M$ have changed, and their respective momentum changes in the same amount and in opposite directions while the whole satisfies the conservation of momentum.
In this section, we will argue the quantumlike nature of gravity based on two no-go theorems we proposed. Before that, we introduce the assumption of locality. Locality is defined here as any interaction between objects at a distance that is mediated through a field, rather than being instantaneous. This asserts that the gravitational interaction between objects $A$ and $B$ is mediated through a gravitational field. That is, there is no direct interaction between objects $A$ and $B$.
Besides, here we would like to emphasize that we assume the gravitational field sourced by a mass [53, 19, 54] is an object like an electromagnetic field. Here we are interested in exploring whether those gravitational fields produced by mass are classical or quantum in nature. Without loss of generality, in the following, we describe the Newtonian gravitational field sourced by matter in terms of a system.
Now we denote the gravitational field labeled as $E$, which is primarily generated by a massive object denoted as $M$. As depicted in Figure 2 (a), at $T=t_{0}$, we suppose there are two stationary quantum masses labeled as $S$ and $M$ respectively. Based on our existing observations of the universe, these masses may mutually attract to each other as a consequence of the gravitational field $E$ mediating their interaction. While the total momentum of the system remains conserved at zero, the individual momentum of system $S$ and system $M$ increases in opposite directions, as shown in Figure 2 (b).
Let us analyze how the predictions of the classical gravitational field contradict the above description of the phenomenon. Here we delve into the discussion of the gravitational field in two cases. In the first case, the gravitational field $E$ is considered to be classical. Due to the local nature of the interaction, the interaction survives only in $ES$ and $EM$. Based on the no-go theorem 1, and assuming momentum conservation, the gravitational field of classical nature $E$ cannot induce any change in the momentum $P$ of quantum matter $S$. In other words,
$\langle P_{S}\rangle_{t}=\langle P_{S}\rangle_{t+\Delta t},$ (11)
where $\langle P_{S}\rangle_{t}=\text{Tr}(P\varrho)$ and $\langle P_{S}\rangle_{t}=\text{Tr}(P\varrho^{\prime})$. It is important to note that here one might consider this result counterintuitive since empirically, the classical gravitational field is capable of changing the momentum of matter. Our conclusion stems from our assumption of momentum conservation for any interacting system. In classical theory, momentum is conserved between mass $S$ and mass $M$ without considering the mediating effect of the gravitational field. Here, the gravitational field is concretized as $E$, and the law of conservation of momentum is assumed to hold during the interaction of $SE$ and $EM$.
Moving on to the second case, let’s now consider $E$ as a quantum gravitational field. In this scenario, the Hamiltonian of the composed system can be expressed as $H_{SME}=H_{S}+H_{E}+H_{M}+H_{SE}+H_{EM}$. Remarkably, the Hamiltonian allows us to meet the momentum conservation criteria, ensuring that $\langle[U_{SEM},P_{S}+P_{E}+P_{M}]\rangle=0$ while $\langle[U_{SEM},P_{S/E/M}]\rangle\neq 0$, where $U_{SEM}=\text{exp}(-iH_{SEM}T/\hbar)$. That is, one may have
$\langle P_{S}\rangle_{t}\neq\langle P_{S}\rangle_{t+\Delta t}.$ (12)
Compared to the classical gravitational field, the quantum one has ability to induce changes in the momentum of a quantum system for both $S$ and $M$. Consequently, if a particle is observed to change momentum after being subjected to a gravitational field, then the gravitational field must be quantum.
Let us now suppose $M$ to be a huge massive object, i.e., Earth or a metal sphere, representing quantum matter that adheres to the classical limit, as shown in Figure 2. According to our argument, under the principle of momentum conservation, any momentum change in the quantum mass $S$ resulting from the influence of $M$ mediated through the gravitational field is viewed as sufficient evidence for the quantum nature of gravity. Actually, there are lots of experimental observations to support our argument. For example, both the free fall of a quantum particle and the displacement of a particle caused by gravitational interaction of the metal sphere [51] offer compelling indications of the quantumness of gravity.
IV Other conservation laws
While we have argued for the non-classical nature of gravity in terms of conservation of momentum, a natural question is why conservation of momentum is special. Is it possible to derive the same conclusion using other conservation laws, such as conservation of energy? The answer is yes.
When we consider the gravitational field as an interacting object, the energy of the particle may be just its own free energy and may not include the gravitational potential energy. Imagine a particle’s energy, defined as free energy, i. e. kinetic energy. Our above analysis asserts that if the kinetic energy of a particle increases following an interaction with a gravitational field, this significant observation implies that gravity behaves in a manner inconsistent with classical objects, assuming the principle of energy conservation holds true [52]. Evidently, the free fall of a particle may provide compelling evidence for the quantum nature of gravity.
V Conclusions
We have discussed a general framework for quantum-classical hybrid dynamics, offering a model-independent approach. The framework is predicated on the fundamental premise that any classical system can only have one basis and maintain classical correlations with a quantum system.
Within the framework, we present a no-go theorem that pertains to systems subject to a conservation law. The no-go theorems explore a fundamental limitation of the classical system to interact with a quantum system, assuming that the interaction satisfies the decomposition property (including Hamiltonian formalism) set out in Figure 1, i. e. any classical system can not induce a change of the local observable of quantum system while a global conserved quantity of hybrid system is assumed. Based on these no-go theories, our exploration delves into the quantum nature of gravity, unveiling a profound connection that exists between conservation laws and the inherent quantum aspects of gravity.
Additionally, our analysis underscores the pivotal role played by expectation values in facilitating a holistic comprehension of the physical universe. This transcends the limitations imposed by disparate mathematical structures and serves as a bridge, connecting various physical theories into a unified description.
While our argument regarding the quantum nature of gravity remains within a non-relativistic context, its generality and validity persist. We expect to discover more relationships between fundamental principles and gravity, leading to a deeper understanding of the nature of spacetime.
Acknowledgements.
T. F. thanks the support of QICI of the University of Hong Kong, and is grateful to Tian Zhang for helpful discussions. C. M. and V.V. thank the Gordon and Betty Moore Foundation. This publication was also made possible through the support of the ID 61466 grant from the John Templeton Foundation, as part of the The Quantum Information Structure of Spacetime (QISS) Project (qiss.fr). The opinions expressed in this publication are those of the Authors and do not necessarily reflect the views of the John Templeton Foundation.